### v2

@@ Abstract
Research surveys have always posed a challenge for beginner researchers who lack of research training. These researchers struggle to understand the directions within their research topic, and the discovery of new research findings within a short time. One way to provide intuitive assistance to beginner researchers is by offering relevant knowledge graphs(KG) and recommending related academic papers. However, existing navigation knowledge graphs primarily rely on keywords in the research field and often fail to present the logical hierarchy among multiple related papers clearly. Moreover, most recommendation systems for academic papers simply rely on high text similarity, which can leave researchers confused as to why a particular article is being recommended. They may lack of grasp important information about the insight connection between ”Issue resolved” and ”Issue finding” that they hope to obtain. To address these issues, this study aims to support research insight surveys for beginner researchers by establishing a hierarchical tree-structured knowledge graph that reflects the inheritance insight of research topics and the relevance insight among the academic papers.
Keywords:
@@
Insight survey Tree-structured Knowledge graph Academic paper Inheritance Relevance
1 Introduction
@@
In recent years, there has been significant progress in the digital environment for academic papers. Databases of academic papers, such as S2ORC[1] and Unarxiv[2], have been continuously improving. The advent of large language models(LLM) like Chatgpt has opened up new possibilities for utilizing these extensive data sources in NLP tasks, including question-answering systems and automatic summarization for academic papers. This advancement enables researchers to conduct research surveys more efficiently using research tools and resources, such as ChatPDF and SciSpace, which have been developed based on ChatGPT and specifically tailored for academic papers understanding[4]. Additionally, knowledge graphs(KG) built on academic papers, such as Connected Papers[3], have also been developed. However, for researchers who are new to a specific academic field or are exploring a research topic, it is difficult to quickly observe, understand, and explore research directions based on large repositories of academic articles[5] solely by simply reading a research paper without a clear target. This inefficient approach to conducting surveys hampers the discovery of research direction and innovative. Therefore, relying solely on paper reading comprehension support software and existing paper recommendation systems may make it difficult to achieve the goal of gaining research insights. Providing an overview of possible research directions and branches within multiple papers on a specific topic may greatly facilitate more efficient exploration of research topics. This process involves the challenging task of conducting in-depth exploration of keywords, theoretical research inheritance, relevance, and their applicability within specific research directions. To address these issue, the objective of this study is to create a knowledge graph for insights survey assistants from multiple academic papers in a specific research topic. It presents a tree structure that shows (1) From the origin of the research task, expand the citation inheritance associations of the research task. (2) Explore the relevance chain among similar research tasks to show relevant research points. It provides potential valuable assistance to researchers in understanding, exploring, and gaining insights into research topics. Additionally, it guides researchers in quickly identifying topic clues and directions for further research expansion. This study contributes to a thorough exploration and visualization of academic topics in the following ways:
@@
(1)Based on the existing academic paper repository, we performed secondary development to fit the requirements of the research insight survey. We integrated a dataset that includes higher-dimensional features, which reflect the citation structure among papers and enable insights content segmentation. (2) To explore the connections of relevance between pairs of papers, we utilize machine learning technology to build a training model for classifying the insight content into ”Issue finding” and ”Issue resolved”. (3) Based on the (1) and (2), we establish a knowledge graph in a hierarchical tree structure. This knowledge graph serves as a guide for beginner researchers exploring research topics to achieve efficient insight survey.
2 Related work
@@
Research survey is a comprehensive investigation conducted to gain an overview of a research field or topic. Especially, it is important for novice researchers to gain a comprehensive understanding of the knowledge structure in their field through research surveys. The development of a KG can be helpful for novice researchers in efficiently understanding and exploring research topics. As previous studies on the KG generated from scientific papers, The PubMed knowledge graph includes meta information such as bio-entities and authors. However, for novice researchers, relying solely on meta information is not sufficient to provide them with enough information about the research content in that field[6]. The knowledge graphs that involve internal information of academic papers include Chan et.al. proposed representation ontology for a four-space integrated KG (background, objectives, solutions, and findings) using NLP technology, as well as Tu et.al.proposed Semantic KG (SKG) that integrates semantic concepts to represent the corpus[7][8].They both classify the content of papers into entity concepts, and associated research key-points in their knowledge graphs. However, the association paths of complex concept maps can be cumbersome, making it difficult for novice researchers to accurately locate academic papers in a specific research direction. This may result in limited insight into the research branch. Jelodar et.al. proposed a semantic framework using probabilistic topic modeling and Gibbs sampling algorithm for a scholarly article recommendation system[9]. However, due to the lack of similarity analysis involving relevance, this makes it difficult for researchers to discover existing issues in a certain research branch. This study aims to expand knowledge mining in the field of insight survey support. It aims to enable researchers to quickly gain insights into research directions. To achieve this, we propose a hierarchical tree structure KG based on the content of academic papers. This graph reflects the inheritance relationships and relevance chain among academic papers.
3 Methodology
3.1 Definition
Research topic: This study focuses on a research topic derived from keywords related to a specific research field. In a research topic, there are multiple research tasks that make up multiple research directions.
Inheritance insight: A survey method that utilizes direct citation relationships between academic papers can be used to identify various branches for exploring inherited relationships in research directions. This study establishes an Inheritance tree to support this kind of insight.
Relevance insight: By analyzing the ”Issue finding” and ”Issue resolved” for each academic paper, a relevance chain of Finding → Resolved → Finding can be established between multiple academic papers to guide researchers in discovering the value of potential direction. This study introduces a relevance tree to facilitate this type of analysis.
@@ 3.1 Definition
The Ideal KG reflecting Inheritance insight and relevance insight survey of ’HotpotQA’ topic are shown in Fig.1-2. The configuration of the KG are shown in Table 1. In Fig.1, the branches p1(Dataset)->p2(Reading Comprehension)->
@@
p4(Evaluation) and p1(Dataset)->p3(Benchmark)->p7(Retrieval) represent two research directions extended by the core paper of the ’HotpotQA dataset’. This graph can help researchers infer the elements of research inheritance in the ’HotpotQA’ topic. For instance, P1 introduced the ’HotpotQA dataset’, P2 utilized this dataset for reading comprehension, and P4 developed evaluation metrics based on reading comprehension. In Figure 2, the path of <Multi-hop branch> p1->Network->p2->Single-hop->p4 and the path of <Retrieval> branch p1->Retrieval->p3->Iteratively->p7 are extensions of the Multi-hop and Retrieval research tasks in the ’HotpotQA’ topic. This tree structure can help researchers infer the correlation factors between subtasks in the ’HotpotQA’ topic. For example, in the ’Multi-hop’ subtask, survey paper P1 provides a list of specific tasks, P2 extends it to the network establishment level, and P3 incorporates single-hop methods.
Table 1: Knowledge graph configuration
@@
Node_label Paper title
Visually Presenting the theme
of a Research Paper
Node_Title
1. Keyword of paper
2. Issue Resolved in this paper
3. Issue Finding of this paper
Visually display the key
information of research paper,
@@ 3. Issue Finding of this paper
Edge_direction
1. Inheritance tree
@@ 2. Relevance tree
1. Inheritance tree
Reflects the Citation direction
2. Relevance tree
Direction of Relevance chain
Edge_label Co-occurring vocabulary
Helps users understand the
relevance between papers.
Edge_Title
The value of the relevance chain
between papers.
Make it convenient for users
to explore potential directions.
Figure 1: Inheritance tree plan(’HotpotQA topic’)
Figure 2: Relevance tree plan(’HotpotQA topic’)
3.2 Implementation procedure
The overview is shown in Fig.3. The development process is divided into 4 sub-stages: 1. We utilize S2orc as our infrastructure data of KG. S2orc is an academic database that encompasses metadata, bibliographic references, and full text for 8.1 million open access papers[1]. 1. In the first stage, we perform secondary processing based on the S2ORC database to build our insight survey dataset. 2. In the second stage, we use the Sentence Bert model and manually set labels to perform three-class classification (Resolved/Neutral/Finding) on the insight content of each paper to extract the corresponding sentence.3. In the third stage, we use the extracted sentences from phase 2 to perform in-depth analysis on inheritance and relevance chain. We select appropriate papers from the whole insight survey dataset based on certain criteria to generate hierarchical trees.4. In the final stage, we use the pyvis[15] to visualize the tree-structured graph and provide some case studies to demonstrate our findings.
Figure 3: Overview
3.3 Phase1 : Data processing
To build the insight survey dataset, the first step is to select a topic and filter out the relevant sub-dataset from the S2orc database. In this study, we will use ’HotpotQA’ as an example.’HotpotQA’, which is a well-known dataset in the field of NLP. We iterate through each paper in the S2ORC database and extract those that contain the keyword ’HotpotQA’ into our sub-paper library. Our sub-set also includes data-citation information, full-text content, and meta-information of the papers associated with ’HotpotQA’. Next, we extract the text from sections titled ’conclusion’ and ’limitation’ to identify insightful content. We focus on paragraphs with section titles containing ’conclusion’ and ’limitation’ as they address the problems discussed in the research paper and highlight any remaining challenges or limitations. Additionally, we extract relevant citation relationships from the S2orc dataset to create a global network based on the annotations of ’bib-entry’ in the S2orc dataset. The details of insight survey content as described in Table 2.
Table 2: Insight dataset processing (HotpotQA)
Annotation
in S2orc
Description Way of Processing
title Paper title
Use the offset annotation of the paper title
to extract the title text.
@@ 3.3 Phase1 : Data processing
To extract paragraphs related to the viewpoints
of ’Finding’ and ’Resolved’ issues, locate a section
header that includes the words ’Conclusion’,
’Discuss’, ’and ’Limitation’. This ensure that the
starting point of the extracted paragraph is appropriate.
@@
Find the corpusid of a paper in S2ORC that matches
the cited paper’s corpusid in the ’HotpotQA’ subset.
3.4 Phase2 : Insight Sentence Extraction
In this section, we discuss the process of insight survey dataset. First, we divide the text in the ’insight-content’ into sentences. We annotate each sentence with the label ”Resolved,” ”Neutral,” or ”Finding” corresponding to the viewpoints the sentence expresses. Next, we use the sentence BERT[10] to vectorize each sentence and adopt SVM classifier to distinguish the corresponding label for each sentence. Finally, sentences with the same ”Resolved” and ”Finding” labels in each article will form the insight sentence of that article.
Sentence Segmentation: Spacy is an open-source natural language processing library for Python that offers an API to access its machine learning trained methods and properties[11]. This work uses the spacy pre-trained model to implement sentence segmentation. Spacy provides a pre-trained English library called en_core_sci_lg which includes a default sentence segmenter [^1] https://allenai.github.io/scispacy/ [^2] https://www.tutorialspoint.com/perform-sentence-segmentation-using-python-spacy. Any complex segmentation patterns that failed were manually fixed.
Human Annotation: For the segmented sentences mentioned above, experts determine the viewpoint of each sentence based on its meaning. The viewpoints include ’Issue Resolved,’ ’Neutral,’ or ’Issue Finding.’ Sentences that thoroughly analyze research methods without explicitly highlighting contributions will be considered ‘Neutral.’ Sentences discussing potential future works in a particular field without specificity will also fall into the ‘Neutral’ category. ’Issue Resolved‘ sentences need at least a combination of contribution, research details, and experimental outcomes, regardless of their positive or negative outcomes. Sentences that ambiguously hint at trends and recommendations may be classified as ‘Issue Finding.’ When categorizing these three viewpoints, the references, tables, and figures of the sentences are unchanged. Table 2 shows specific definitions, distinguishing criteria, and examples of these three labels. The dataset, which consists of insight sentences and labels, has been published on  [^3] https://www.kaggle.com/datasets/dannyleeakira/dataset-for-academic-novelty-insight-survey
Table 3: Sample - Human annotation strategy
Sentence Description Label
In this paper, we present textbrewer, a flexible pytorch-
based distillation toolkit for nlp research and applications.
The target of
this paper
Resolved
@@ 3.4 Phase2 : Insight Sentence Extraction
by learning an explicit hierarchy…
Things achieved
in this paper
Resolved
@@
on the fever dataset, followed by 4 epochs on liar-plus.
Detail of the
method they did
Neutral
We have conducted a series of experiments.
Not related to the
Insight viewpoints
Neutral
@@
downstream nlp applications where the graph…
Future work of
this paper
Finding
@@
Limitation of
this paper
Finding
Training, Sentence Extraction and evaluation: We adpot ’Distilbert-base-uncased’ pre-training model, which was trained using Wikipedia and BookCorpus, for sentence vectorization. We selected Support Vector Machines (SVM) for classifying the vectorized sentences due to their strong generalization performance. We also used 947 labeled sentences for training and validation data. The details of the training data and test data are shown in Table 5. To obtain the optimal SVM parameters, we use the grid search[12] method and find the best parameters to apply to the test. The classification accuracy evaluation is presented in Table 5. There is an imbalance in the number of instances across classes. ”Resolved” class has significantly more instances than the other classes in both the training and test sets. Classification result also shows that ”Resolved” class has higher F1-score as it might be influenced by the larger amount of data. While the ”Neutral” and ”Finding” class has lower F1-scores, indicating challenges in achieving both high precision and recall. Based on the results, we extracted insight sentences of the ”Resolved” and ”Finding” by combining the sentences with corresponding labels within each article.
Table 4: Detail of issue-status dataset
Table 5: Classification result of issue status
@@
Resolved 438 142
Neutral 146 53
Finding 126 42
Total 710 237
@@
Resolved 0.97 0.81 0.88
Neutral 0.40 0.61 0.48
Finding 0.53 0.68 0.60
Table 5: Classification result of issue status
3.5 Phase3-4 : Hierarchical Tree Construction & Visualization
In this section, we utilize the sentences extracted in phase 2 and the citation information obtained in phase 1 to comprehensively extract the characteristics of the paper. We then employ two strategies to select specific papers in order to construct a tree-structured network.
Similarity calculation: To create the relevance tree, we use the classification results from Phase 2 to determine the elements of the relevance chain (Issue finding → Issue Resolved). We then use embedded insight sentences in sentence transformers[14] to calculate the cosine similarity between insight sentences labeled as ’Finding’ and those labeled as ’Resolved’. Next, we iterate through all the insight sentences in the papers and calculate the relevance chain to generate the relevance matrix. Based on the values in the relevance matrix, we select papers to construct the relevance tree in the following section.

### v5

@@ Abstract
Research surveys have always posed a challenge for novice researchers who lack research training. These researchers struggle to understand the directions within their research topic and the discovery of new research findings within a short time. One way to provide intuitive assistance to novice researchers is by offering relevant knowledge graphs (KG) and recommending related academic papers. However, existing navigation knowledge graphs mainly rely on keywords or meta information in the research field to guide researchers, which makes it difficult to clearly present the hierarchical relationships, such as inheritance and relevance between multiple related papers. Moreover, most recommendation systems for academic papers simply rely on high text similarity, confusing researchers as to why a particular article is recommended. They may lack the grasp of important information about the insight connection between ‘Issue resolved’ and ‘Issue finding’ that they hope to obtain. This study aims to support research insight surveys for novice researchers by establishing a hierarchical tree-structured knowledge graph that reflects the inheritance insight and the relevance insight among multiple academic papers on specific research topics to address these issues.
Index Terms:
@@
Insight survey, Tree-structured, Knowledge graph, Academic paper, Inheritance, Relevance, Research direction, Academic issue;
I Introduction
@@
In recent years, there has been significant progress in the digital environment for academic papers. Databases of academic papers, such as S2orc[1] and Unarxiv[2], have been continuously improving. The advent of large language models (LLM) like Chatgpt have opened up new possibilities for utilizing these extensive data sources in NLP tasks, including question-answering systems and automatic summarization for academic papers. This advancement enables researchers to conduct research surveys more efficiently using research tools and resources, such as ChatPDF and SciSpace, which have been developed based on ChatGPT and specifically tailored for academic papers understanding[4]. Additionally, knowledge graphs (KG) built on academic papers, such as Connected Papers[3], make it more convenient for researchers to locate desired papers through keyword search. However, for researchers who are new to an academic field or who are exploring a specific research topic, it is difficult to quickly observe, understand, and explore research directions based on large repositories of academic articles solely by simply reading academic papers without a clear target[5]. This inefficient survey approach hampers the discovery of research direction and innovation. Therefore, relying solely on paper reading comprehension support software and existing paper recommendation systems may make it difficult to achieve the goal of gaining research insights. Providing an overview of possible research directions and branches within multiple papers on a specific topic may greatly facilitate a more efficient exploration of research topics. This process involves the challenging task of conducting an in-depth exploration of keywords, inheritance, relevance, and their applicability within specific research directions. To address these issues, the objective of this study is to create a knowledge graph for insights survey assistants from multiple academic papers on a specific research topic. It presents a tree structure that shows:
(1) From the origin of the research task, expand the citation inheritance and relevance associations . Here, ‘Inheritance’ refers to the route of research evolution from the origin (history) of a research direction to the current stage, which consists of at least two or more academic papers. ‘Relevance’ refers to the path composed of multiple academic papers that aim to solve similar academic issues, which is formed by the relevance chain. Unlike ‘Inheritance’ depended on citation connection, relevance tends to link the similarity of existing academic issues to discover the research direction.
(2) Explore the relevance chain within similar research tasks to demonstrate relevant research points. It provides potential valuable assistance to researchers in understanding, exploring, and gaining insights into research topics. Additionally, it guides researchers in quickly identifying topic clues and directions for further research expansion. This study contributes to an in-depth exploration and visualization of research topics in the following ways:
@@
1. Based on the existing academic paper repository, we performed secondary development to fit the requirements of the research insight survey. We integrate a dataset that includes higher-dimensional features, which reflect the citation structure among papers and enable insights content segmentation.
2. To explore the connections of academic issue relevance between pairs of papers, we utilize machine learning technology to build a training model for classifying the insight content into ‘Issue finding’ and ‘Issue resolved’.
3. Based on the (1) and 2. , we establish a KG in a hierarchical tree structure. This KG serves as a guide for novice researchers exploring research topics to achieve efficient insight surveys.
II Related work
@@
Research survey is a comprehensive investigation conducted to gain an overview of a research field or topic. Novice researchers need to comprehensively understand the knowledge structure in their field through research surveys. The development of KG can benefit for novice researchers in efficiently comprehending and exploring research topics. As previous studies on the KG generated from academic papers, Xu et al. constructed a PubMed knowledge graph that includes meta information such as bio-entities and authors[6] and Martha et al. proposed the Tree of Science (ToS), which recommends articles based on their position in the graph of citation[7]. However, for novice researchers, relying solely on meta or citation information is insufficient to provide enough information about the research content in that field to gain the overview concept map of the research branch. Deeper data mining from the academic paper is still required to refine the academic knowledge graph, which may provide high-quality guidance for the research survey. The knowledge graphs that involve internal information of academic papers include Chan et al. proposed representation ontology for a four-space integrated KG (background, objectives, solutions, and findings) using NLP technology, as well as Tu et al. proposed Semantic Knowledge Graph (SKG) that integrates semantic concepts to represent the corpus[8][9]. They both classify the content of papers into entity concepts and associated research key-points in their knowledge graphs. However, the association paths of complex concept maps can be cumbersome, making it difficult for novice researchers to accurately locate academic papers in a specific research direction, which may result in limited insight into the research branch. Thus, extracting insightful content from the paper is crucial to guide researchers in conducting effective surveys. Hayashi et al. introduce disentangled paper summarization, where models generate both contribution and context summaries simultaneously to hint at the contribution overview of a specific research direction[10]. This type of summary may also link the contributions of multiple papers to perform a research survey of novelty in that research topic. However, the absence of summaries related to ‘issues that could be improved’ makes it challenging for researchers to trace the evolution history in a particular research branch, such as the relevance of ‘existing issues’ and ‘issues need to be resolved.’ The above studies have yet to involve the intuitive and concise expression of the evolution trajectory and the implicit connection between multiple papers. Therefore, We are committed to mining the inheritance relationships among papers and exploring the relevant connections between their mentioned issues. We focus on expanding knowledge mining in the field of insight survey support, which aims to enable researchers to gain insights into research directions quickly. To achieve this, we propose a hierarchical tree-structured KG based on the insight content of academic papers that reflects the inheritance relationships and relevance chain among academic papers.
III Methodology
Figure 1: Inheritance tree plan(‘HotpotQA topic’)
Figure 2: Relevance tree plan(‘HotpotQA topic’)
@@ III Methodology
In this work, we will use ‘HotpotQA’ as an example, which is a well-known dataset in the field of NLP[11]. The Ideal KG reflecting Inheritance insight and relevance insight survey of ‘HotpotQA’ topic are shown in Fig.1-2. The configuration of the KG is shown in Table I. In Fig.1, the branches:
p1(Dataset)-> p2(Reading Comprehension)
-> p4(Evaluation) and
@@
p1(Dataset)-> p3(Benchmark)-> p7(Retrieval) represent two research directions extended by the core paper of the ‘HotpotQA dataset’. This graph can help researchers infer the elements of research inheritance in the ‘HotpotQA’ topic. For instance, P1 introduced the ‘HotpotQA dataset’, P2 utilized this dataset for reading comprehension, and P4 developed evaluation metrics based on reading comprehension. In Figure 2, the path of ‘Multi-hop’ branch:
p1 -> Network -> p2 -> Single hop -> p4
and the path of ‘Retrieval’ branch:
p1 -> Retrieval -> p3 -> Iteratively -> p7
are extensions of the ‘Multi-hop’ and ‘Retrieval’ research tasks in the ‘HotpotQA’ topic. This tree structure included multiple paths can help researchers infer the correlation factors between subtasks in the ‘HotpotQA’ topic. For example, in the ‘Multi-hop’ subtask, survey paper P1 provides a list of specific tasks, P2 extends it to the network establishment level, and P3 incorporates single-hop methods.
TABLE I: Knowledge graph configuration
@@
Node__\__label Paper title Visually Presenting the theme of a Research Paper
Node__\__Title
1. Keywords of paper
2. ‘Issue Resolved’ in this paper
3. ‘Issue Finding’ in this paper
Visually display the key information of research paper,
@@ 1. Keywords of paper
Edge__\__direction
1. Inheritance tree :
@@ 2. Relevance tree :
1. Inheritance tree : Reflects the citation direction
2. Relevance tree : Direction of Relevance chain
Edge__\__label Co-occurring vocabulary
Helps user understanding of the
relevance between the two papers.
Edge__\__Title The value of relevance chain between papers Make it convenient for users to explore potential direction
III-A Implementation procedure
We utilize S2orc as our infrastructure data of KG. S2orc is an academic database encompassing metadata, bibliographic references, and full text for 8.1 million open access papers[1]. The overview of implementation is shown in Fig.3. The development process is divided into 4 sub-stages:
1. In the first stage, we meticulously manipulate secondary data based on the S2orc database to construct our unique dataset tailored specifically for our insight survey.
2. In the second stage, we use the Sentence Bert model and manually set labels to perform three-class classification (Issue Resolved/Neutral/Issue Finding) on the insight content of each paper to extract the corresponding sentence.
3. In the third stage, we use the extracted sentences from phase 2 to analyze the inheritance and relevance chain in-depth. We select appropriate papers from the whole insight survey dataset based on certain criteria to generate tree-structured hierarchical trees.
4. In the final stage, we use the pyvis[19] to visualize the tree-structured KG and provide some case studies to demonstrate our findings.
Figure 3: Implementation procedure
III-B Phase1 : Data processing
To build the insight survey dataset, the first step is to select a topic and filter out the relevant sub-dataset from the S2orc dataset. We iterate through each paper in the S2orc dataset and extract those that contain the keyword ‘HotpotQA’ into our sub-set. Our sub-set also includes data-citation information, full-text content, and meta-information of the papers associated with ‘HotpotQA.’ Next, we extract the text from sections titled ‘conclusion’,‘discussion’ and ‘limitation’ to identify insight content. We focus on paragraphs with section titles containing ‘conclusion’, ‘discussion’ and ‘limitation’ as they address the problems discussed in the research paper and highlight any remaining challenges or limitations. Additionally, we extract relevant citation relationships from the S2orc dataset to create a global citation network based on the annotations of ‘bib-entry’ . The details of insight survey content as described in Table II.
TABLE II: Insight dataset processing (HotpotQA)
Annotation in S2orc Description Way of Processing
title Paper title Use the offset annotation of the paper title to extract the title text.
@@ III-B Phase1 : Data processing
To extract paragraphs related to the viewpoints of ‘Finding’ and
’Resolved’ issues, locate a section header that includes the words
‘Conclusion’, ‘Discuss’, and ‘Limitation’. This ensure that the starting
point of the extracted paragraph is appropriate.
@@
Find the corpusid of a paper in S2orc that matches the cited paper’s
corpusid in the ‘HotpotQA’ subset.
III-C Phase2 : Insight Sentence Extraction
In this section, we discuss the process of the insight survey dataset. First, we divide the text in the ‘insight-content’ into sentences. We annotate each sentence with the label ‘Issue Resolved’,‘Neutral’, or ‘Issue Finding’ corresponding to the viewpoints the sentence expresses. Next, we use the Sentence Bert[12] to vectorize each sentence and adopt Support Vector Machines (SVM) classifier to distinguish the corresponding label for each sentence. Finally, sentences with the same ‘Issue Resolved’ and ‘Issue Finding’ labels in each article form the ‘insight sentence’ of that article.
Sentence Segmentation: Spacy is an open-source natural language processing library for Python that offers an API to access its machine learning trained methods and properties[13]. This work uses the pre-trained model in Spacy to implement sentence segmentation. Spacy provides a pre-trained English library called ‘en__\__core__\__sci__\__lg’ which includes a default sentence segmenter [^1] https://allenai.github.io/scispacy/ [^2] https://www.tutorialspoint.com/perform-sentence-segmentation-using-python-spacy. Any complex segmentation patterns that failed were manually fixed.
Human Annotation: For the segmented sentences mentioned above, experts determine the viewpoint of each sentence based on its meaning. The viewpoints include ’Issue Resolved,’ ’Neutral,’ or ’Issue Finding.’ Sentences that thoroughly analyze research methods without explicitly highlighting contributions will be considered ‘Neutral.’ Sentences discussing potential future works in a particular field without specificity will also fall into the ‘Neutral’ category. ’Issue Resolved‘ sentences need at least a combination of contribution and experimental outcomes, regardless of their positive or negative outcomes. Sentences that ambiguously hint at trends and recommendations may be classified as ‘Issue Finding.’ When categorizing these three viewpoints, the references, tables, and figures of the sentences are unchanged. Table 2 shows specific definitions, distinguishing criteria, and examples of these three labels. The dataset, which consists of insight sentences and labels, has been published on  [^3] https://www.kaggle.com/datasets/dannyleeakira/dataset-for-academic-novelty-insight-survey
TABLE III: Human annotation strategy
Sentence Description Label(Issue)
In this paper, we present textbrewer, a flexible pytorch-based
distillation toolkit for nlp research and applications.
The target of this paper Resolved
@@ III-C Phase2 : Insight Sentence Extraction
by learning an explicit hierarchy of tasks based on the skills
they require.
Things achieved in this paper Resolved
@@
on the fever dataset, followed by 4 epochs on liar-plus,
the fine-tune on politihop for 4 epochs.
More detail of the method they did Neutral
We have conducted a series of experiments. Not related to the insight viewpoints Neutral
@@
downstream nlp applications where the graph inductive
bias is necessary and dataset is scarce.
Future work of this paper Finding
@@
Limitation of this paper Finding
Training, Sentence Extraction and evaluation: We adopt ‘scibert__\__scivocab__\__uncased’ pre-training model [^4] https://huggingface.co/allenai/scibert__\__scivocab__\__uncased [14], which was trained comprising 1.14M full-papers and 3.1B tokens, was sourced from Semantic Scholar, for sentence vectorization. scibert__\__scivocab__\__cased exhibits adaptability to both the corpus and domain, making it suitable for our training data. We selected SVM for classifying the vectorized sentences due to their strong generalization performance. We also used 1500 labeled sentences for training and validation data. The training and test data details are shown in Table IV. To obtain the optimal SVM parameters, we use the grid search[15] method and find the best parameters to apply to the test. The classification accuracy evaluation is presented in Table V. The classification result shows that the ‘Issue Resolved’ class has a higher F1 Score, as the larger amount of data might influence it. The ’Neutral’ and ’Issue Finding’ classes have lower F1 scores, indicating challenges in achieving both high precision and recall. Based on the results, we extracted insight sentences of the ‘Issue Resolved’ and ‘Issue Finding’ by combining the sentences with corresponding labels within each article.
TABLE IV: Detail of issue-status dataset
@@
Issue Resolved 532 165
Neutral 334 121
Issue Finding 259 89
Total 1125 375
TABLE V: Classification result of issue status
@@
Issue Resolved 0.90 0.85 0.88
Neutral 0.62 0.73 0.67
Issue Finding 0.75 0.71 0.73
III-D Phase3-4 : Hierarchical Tree Construction & Visualization
In this section, we utilize the sentences extracted in phase 2 and the citation information obtained in phase 1 to comprehensively extract the insight characteristics of the papers. We then employ two strategies to select specific papers to construct a tree-structured network.
Similarity calculation: To create the relevance tree, we use the classification results from Phase 2 to determine the elements of the relevance chain (‘Issue finding’ → ‘Issue Resolved’). We then use embedded insight sentences in sentence transformers[18] with ‘scibert__\__scivocab__\__uncased’ model to calculate the cosine similarity between insight sentences labeled as ‘Issue Finding’ and those labeled as ‘Issue Resolved.’ Next, we iterate through all the insight sentences in the papers and calculate the relevance chain to generate the relevance matrix. Based on the values in the relevance matrix, we select papers to construct the relevance tree in the following section.